# Principles for Optimizing Quantum Transduction in Piezo-Optomechanical Systems

### v2

@@ IV.2.2 Analysis of optomechanical coupling
The most straightforward way to change to the effective optomechanical coupling $\bar{G}\bar{a}$ is to change the power of the pump laser, which changes $\bar{a}$. When this is not reasonable (e.g., it approaches optical damage thresholds), we can see how to change $\bar{G}$ itself both through bulk material properties and design.
@@
Regarding bulk material properties, we see from (IV.2.1) that $\bar{G}_{mn}$ scales linearly with the dielectric constant $\epsilon_{ij}$, multiplied by the photoelasticity tensor $p_{ijk\ell}$, and divided by the square root of the density of the material. In Appendix A, we have tabulated the properties of common optical materials, and one can see for example that barium titanate is a promising candidate for a stronger optomechanical coupling than the silicon nitride used in the MRRs in [16]. Crystalline silicon also has a higher optomechanical figure of merit than silicon nitride, but its overall optomechanical coupling is limited due to the two-photon absorption that occurs in silicon at higher powers due to its large third-order optical nonlinearity.
For aspects of the optomechanical coupling that can be effected by design, we must look to optimize the overlap integral in $\bar{G}_{mn}$ (IV.2.1). Because the overlap integral has the product of two electric fields instead of one multiplied with the strain, we maximize the overlap not when the wavelengths of the sound and light are equal, but when the wavelength of the sound is approximately half of the wavelength of the light. This can be attributed to an optomechanical analogue of what is known as phase matching in nonlinear optics, ensuring conservation of momentum between the photons and phonons in the transduction process at their respective frequencies  [^5] In order to show this explicitly, one would have to convert the complete transduction hamiltonian into the basis of the optical symmetric and asymmetric supermodes, but it is more straightforward to compute the transduction efficiency as shown in the rest of this paper.. At GHz-scale frequencies, this may be achievable, albeit with some challenge in tuning the optical frequency to the required precision. Alternatively, one can have the sound oscillate uniformly over the dimensions of the MRR by having it propagate parallel to its axis of symmetry (perpendicular to the direction of light propagation). This is the design considered in this setup, which also affords a more broadband coupling interaction. Moreover, this transverse coupling scheme does not violate phase matching/momentum conservation, as the thickness of the MRR in this transverse direction is so small that it is generally less than the wavelength of the acoustic phonon at these frequencies. Because of this, there are a broad range of transverse acoustic phonons that could be optomechanically phase-matched to the photons propagating in the MRR, and it is not difficult to simultaneously satisfy both energy and momentum conservation for this interaction.
IV.3 The optical (evanescent) coupling constants
@@
One special aspect of the design of this transduction platform is the use of a pair of coupled ring resonators, and using their symmetric and asymmetric supermodes as the principal optical resonances. A simpler design would be to use just one micro-ring resonator, optomechanically coupled to the piezoelectric medium, and evanescently coupled to the optical bus waveguide. Such a simplified system can in principal have transduction efficiencies comparable to what is seen in the more sophisticated double-ring platform studied here, but at a great cost that compromises its utility.
@@
The resonances of a single MRR are (nominally) equally spaced in frequency, relative to the integer number of wavelengths occurring in a round trip, and the dispersion in the MRR. When a single resonance is excited with a classically bright pump laser to facilitate transduction between adjacent resonances, one also generates light at the transduced frequency due to (nondegenerate) spontaneous four wave mixing (SFWM), where pairs of generated photons at opposite adjacent resonances are created by annihilating pairs of photons at the pump frequency resonance. This can occur in these systems because the sum of the momenta of a photon one free spectral range (FSR, i.e., the spacing between resonances) below the pump frequency and one photon an FSR above the pump frequency is approximately equal to the sum of the momenta of two photons at the pump frequency, satisfying both energy and momentum conservation (which is commonly called being phase-matched). The relatively large number of photon pairs generated by SFWM, compared to the single photons we seek to obtain as a result of transduction would overwhelm any transduction signal with noise. Moreover, with all resonances equally spaced more or less, and the spacing needing to be only a few GHz apart for microwave-optical transduction there would be more light generated at many other pairs of resonances due to their being phase-matched, which at high intensity can also feed back into the transduced mode, contributing still more noise photons at the transduction frequency. On top of this, having an MRR large enough to have an FSR in the GHz scale would have a radius of the order of a centimeter, making integration challenging at a large scale.
@@
By utilizing a pair of MRRs for the optical component of the transducer instead of just one, the frequency resonances are similar to the single-MRR case, except that each single peak is split into two adjacent peaks, whose spacing is independent of the round-trip time in the MRRs (See Fig. 5 for diagrammatic plot). By making the spacing between the split peaks equal to the microwave frequency, one is free to have a much wider FSR, employing much smaller MRRs (and thus a much smaller footprint) which will consequently have much smaller round trip times, higher finesses, and a larger cavity enhancement factor so that smaller pump powers may be used to drive the system at the same intensity as would be needed in the single-ring system. Because we can choose the FSR to be much larger than the frequency splitting, and any light generated by SFWM in peaks equally spaced from the pump resonance would be so far apart that they are either not phase matched due to dispersion (inhibiting the generation of SFWM light) or else the SFWM light that is generated is still readily filtered out from the transduced light due to the large frequency separation. If we pump one of these resonant peaks with classically bright light, its adjacent peak (from the frequency splitting) will not be well-phase-matched for SFWM since there is no corresponding resonance on the opposite side of the pump resonance with the same frequency separation. Thus, with the double MRR design, we nearly eliminate SFWM as a confounding source of noise at the transduction frequency without compromising transduction efficiency. SFWM is not entirely eliminated, but is greatly suppressed relative to the one-ring case where all wavelengths are resonant. This does not completely eliminate other confounding third-order nonlinear-optical effects, such as the pump intensity effectively changing the index of refraction of the material (known as self-phase modulation, of which electrostriction contributes a small part).
@@
Figure 5: Comparison of transmission spectra through the bus waveguide coupled to either a single ring/double-bus (dashed curve) or a double-ring/double bus (solid curve) micro ring resonator system. Here, we see that for two identical rings of round-trip time $T$, that the inter-MRR coupling introduces a splitting of the resonances independent of the free-spectral range of the rings. For this plot, we used critical coupling where the round-trip loss in one MRR is equal to the “reflection” coefficient off of the bus waveguide back into the MRR. Losses and coupling parameters for this plot were chosen to illustrate the structure of the transmission spectrum. By utilizing this splitting, we can use small rings with a large FSR (and therefore large finesse and cavity enhancement factor) whilst keeping the spacing between resonant frequencies of interest in the microwave band to facilitate transduction. Note: if we were to include optical backscattering, this would amount to including a second splitting on top of the already one split peaks (resulting in quadruplets). In assuming the optical backscatter is negligible, we assume this second splitting is too small to be resolved, relative to the linewidth.
IV.3.1 Inter-MRR coupling constant
@@
which is the form describing the system under study here. From this, we may understand that the evanescent optical coupling $J$ between MRRs determines the frequency splitting between the symmetric and antisymmetric supermodes used in this transduction scheme. Tuning this evanescent coupling amounts to changing the spacing between the MRRs, as described in the next subsection.
IV.3.2 waveguide coupling constant
The waveguide coupling constant $\kappa_{ex,2}$ describes the fraction of evanescent light surrounding the bus waveguide that couples into the MRR next to it and vise versa.
To adjust this coupling, one can either change the distance between the bus waveguide and the MRR, or else change the shape of the MRR to increase the interaction distance with the waveguide.
In general, the optical waveguide coupling is dependent on the wavelength of the light being coupled. If we consider two parallel waveguides close enough to one another that the evanescent field surrounding one waveguide overlaps with the other, we can express the light propagating in the waveguides in a new basis. Instead of separate modes interacting with one another, we can consider superpositions of modes (i.e., supermodes) that are independent of one another. Light initially in one waveguide, is in an even superposition of the symmetric and antisymmetic supermodes, but these modes in general have different mode profiles, and as a result, experience different effective indices of refraction $(n_{eff}^{(sym)},n_{eff}^{(asym)})$ throughout the waveguides. Because of this accumulating phase lag between supermodes, the light overall oscillates between waveguides with a characteristic coupling beat length $L_{c}$ [36]:
$L_{c}\approx\frac{\lambda}{2(n_{eff}^{(sym)}-n_{eff}^{(asym)})}$ (67)
@@
In the transducer system considered here, we can understand that if we are considering GHz-scale detunings from a telecom-band wavelength, this is a correspondingly small change in the coupling beat length $L_{c}$, and therefore the coupling constant. For the small detunings considered in this paper, we may take the waveguide optical coupling to be constant, but still something that may be set at the time of manufacture.
V Discussion: Optimizing the transduction efficiency:
Tuning the couplings
@@ V Discussion: Optimizing the transduction efficiency:
Using the equations coupling the different frequency modes in the transducer (III.2.2), one can derive an overall input-output relation for the transducer, as shown in Section III.3:
$\tilde{a}_{out}(\omega)$ $=G_{\tilde{a}_{out},\tilde{c}_{in}}(\omega)\tilde{c}_{in}(\omega)+...$ (68a)
$G_{\tilde{a}_{out},\tilde{c}_{in}}$ $=\frac{\sqrt{\kappa_{ex,2}}\sqrt{\gamma_{ex}}\chi_{01}\chi_{02}\chi_{m}\bar{G}\bar{a}_{1}J}{1+\bar{G}^{2}|\bar{a}_{1}|^{2}\chi_{01}\chi_{m}+J^{2}\chi_{01}\chi_{02}}$ (68b)
@@
Starting with the system parameters in Ref. [16], we can reproduce the efficiency plot as a function of driving laser power, as shown in Fig. 6 (shallow solid blue curve).
To optimize the transduction efficiency, we can tune three key parameters: the optomechanical coupling rate, the electromechanical coupling rate, and the external optical waveguide coupling rate.
It is simplest to tune the optomechanical coupling rate, as this quantity only enters into the transduction efficiency formula via the driving laser amplitude  [^6] This is under the assumption that the optomechanical coupling is not a substantial source of loss via electrostriction contributing to $\kappa_{0.2}$.. Because of this, changing the optomechanical coupling does not change the maximum transduction efficiency, but it does change the required laser power needed in order to achieve it.
Changing the electromechanical coupling $g_{EM}$ simultaneously changes $\gamma_{ex}$ (the coupling of microwave photons in/out of the bulk acoustic resonator) and $\gamma_{m}$ (the overall mechanical damping constant due to both loss and deliberate coupling) while keeping $\gamma_{0}$ (the intrinsic mechanical coupling describing acoustic loss) fixed. Overall, we see that we can increase the maximum transduction efficiency by increasing this coupling, though that maximum appears at higher powers as well (see dotted green curve in Fig. 6).

### v4

@@ III.2.2 Analysis of optomechanical coupling
The most straightforward way to tune the effective optomechanical coupling $\bar{G}\bar{a}$ is to change the power of the pump laser, which changes $\bar{a}$. When this is not reasonable (e.g., it approaches optical damage thresholds), we can see how to change $\bar{G}$ itself both through bulk material properties and design.
@@
Regarding bulk material properties, we see from (III.2.1) that $\bar{G}_{mn}$ scales linearly with the dielectric constant $\epsilon_{ij}$ (owing to its inverse scaling with $\eta_{ij}$), multiplied by the photoelasticity tensor $p_{ijk\ell}$, and divided by the square root of the density of the material. Unlike the piezoelectric interaction, the dielectric constant here is taken at optical frequencies, and is approximately equal to the square of the refractive index. In Appendix A, we have tabulated the properties of common optical materials, and one can see for example that barium titanate is a promising candidate for a stronger optomechanical coupling than the silicon nitride used in the MRRs in [16]. Crystalline silicon also has a higher optomechanical figure of merit than silicon nitride, but its overall optomechanical coupling is limited due to the two-photon absorption that occurs in silicon at higher powers due to its narrow bandgap.
For aspects of the optomechanical coupling that can be affected by design, we must look to optimize the overlap integral in $\bar{G}_{mn}$ (III.2.1). Starting with polarization, we can focus on how the crystal symmetries affect the photoelasticity tensor, and how that may ultimately affect the optomechanical coupling.
$\bar{G}_{mn}\propto\int\!\!d^{3}r\,p_{ijk\ell}^{(mn),eff}\mathcal{E}_{(i,m)}(\vec{r})\mathcal{E}^{*}_{(j,m)}(\vec{r})\frac{\partial w_{(k,n)}(\vec{r})}{\partial r_{\ell}}$ (23)
For each crystal class, the symmetries of the photoelastic tensor $p_{ijk\ell}$ are nearly identical to the corresponding symmetries of the (mechanical) elastic tensor $c_{ijk\ell}$ with the exception that $p_{ijk\ell}\neq p_{k\ell ij}$. This means that while the photoelastic tensor can be written as a $6\times 6$ matrix using Voigt notation, it is not guaranteed to also be a symmetric matrix. Knowing this, the symmetries in the respective upper and lower triangular parts of the photoelastic tensor are inherited individually from the corresponding symmetries in the mechanical elastic tensor. We point this out due to the relative abundance of information on the symmetries of the mechanical elastic tensor.
With the symmetries of the photoelastic tensor in mind, and because we have only one optical pump field, we need only consider components in the form of $p_{ijk\ell}$ where $i=j$ when trying to optimize the optomechanical coupling with regards to polarization. For most of the crystal classes of tabulated materials with significant photoelasticity, we find that the significant elements of the photoelasticity tensor are the ones that connect components of normal strain to components of the electric field. In other words, we generally need to consider only the underlying $3\times 3$ matrix $p_{iijj}$.
Where the piezoelectric and optomechanical media are mechanically bound to one another, the sound waves generated by the piezoelectric medium will dictate what strains need be considered in the optomechanical coupling component. For example, where the electrodes in the transducer are placed to generate an oscillating electric field in the vertical ($z$) direction through the piezoelectric medium, we know that this predominantly produces normal strains $x_{zz}$ also propagating in the $z$ direction. Thus, the elements to the photoelasticity tensor significant to this transducer (in Voigt notation) will be $p_{13}$, $p_{23}$ and $p_{33}$. In Table 4, we have tabulated values of $p_{33}$, which gauge the extent to which this sound can generate light polarized in the same direction (i.e., TM light polarized out of plane), propagating through the resonator. Where it is more common to fabricate resonators that support single TE (polarized in-plane) optical modes, the important elements of the photoelasticity tensor will be $p_{13}$ and $p_{23}$. Note that because the optical field appears twice in the overlap integral, there is no overall cancellation when integrating over the dimensions of the MRR.
Where conventional integrated photonics employ single-mode waveguides for TE light (polarized in the plane of the medium), MRRs made in this design will have to have maximal $p_{13}$ and $p_{23}$ for maximum optomechanical coupling, though they need not both be large for the coupling to be significant.
To maximize the spatial component of the overlap integral, we use arguments similar to those of phase-matching in nonlinear optics. The MRR is very thin in the $z$-direction, even when compared to the wavelength of the GHz phonons propagating through it. Because of this, we need only ensure that the MRR is located within a vibrational antinode to maximize the strain, as well as its overlap with the optical field.
III.3 The optical (evanescent) coupling interaction
@@
One special aspect of the design of this transduction platform is the use of a pair of coupled ring resonators, and using their symmetric and antisymmetric supermodes as the principal optical resonances. A simpler design would be to use just one micro-ring resonator, optomechanically coupled to the piezoelectric medium, and evanescently coupled to the optical bus waveguide. Such a simplified system can in principal have transduction efficiencies comparable to what is seen in the more sophisticated double-ring platform studied here, but at a great cost that compromises its utility.
@@
The resonances of a single MRR are (nominally) equally spaced in frequency, relative to the integer number of wavelengths occurring in a round trip, and the dispersion in the MRR. When a single resonance is excited with a classically bright pump laser to facilitate transduction between adjacent resonances, one also generates light at the transduced frequency due to (nondegenerate) spontaneous four wave mixing (SFWM), where pairs of generated photons at opposite adjacent resonances are created by annihilating pairs of photons at the pump frequency resonance. This can occur in these systems because the sum of the momenta of a photon one free spectral range (FSR, i.e., the spacing between resonances) below the pump frequency and a photon one FSR above the pump frequency is approximately equal to the sum of the momenta of a pair of photons at the pump frequency, satisfying both energy and momentum conservation (which is commonly called being phase-matched). The relatively large number of photon pairs generated by SFWM, compared to the single photons we seek to obtain as a result of transduction would overwhelm any transduction signal with noise. Moreover, with all resonances more or less equally spaced, and the spacing needing to be only a few GHz apart for microwave-optical transduction, there would be more light generated at many other pairs of resonances due to their being phase-matched, which at high intensity can also feed back into the transduced mode, contributing still more noise photons at the transduction frequency. On top of this, having an MRR large enough to have an FSR in the GHz scale would have a radius of the order of centimeters, making integration challenging at a large scale.
@@
By utilizing a pair of MRRs for the optical component of the transducer instead of just one, the frequency resonances are similar to the single-MRR case, except that each single peak is split into two adjacent peaks, whose spacing is independent of the round-trip time in the MRRs (See Fig. 3 for diagrammatic plot). By making the spacing between the split peaks equal to the microwave frequency, one is free to have a much wider FSR, employing much smaller MRRs (and thus a much smaller footprint) which will consequently have much smaller round trip times, higher finesses, and a larger cavity enhancement factor so that smaller pump powers may be used to drive the system at the same intensity as would be needed in the single-ring system. Because we can choose the FSR to be much larger than the frequency splitting, any light generated by SFWM in peaks equally spaced from the pump resonance can be made so far apart that they are either not phase matched due to dispersion (inhibiting the generation of SFWM light) or else the SFWM light that is generated is still readily filtered out from the transduced light due to the large frequency separation. If we pump one of these resonant peaks with classically bright light, its adjacent peak (from the frequency splitting) will not be well-phase-matched for SFWM since there is no corresponding resonance on the opposite side of the pump resonance with the same frequency separation. Thus, with the double MRR design, one nearly eliminates SFWM as a confounding source of noise at the transduction frequency without compromising transduction efficiency. SFWM is not entirely eliminated, but is greatly suppressed relative to the one-ring case where all wavelengths are resonant. This does not completely eliminate other confounding third-order nonlinear-optical effects, such as the pump intensity effectively changing the index of refraction of the material (known as self-phase modulation, of which electrostriction contributes a small part).
@@
Figure 3: Comparison of transmission spectra through the bus waveguide coupled to either a single ring/double-bus (dashed curve) or a double-ring/double bus (solid curve) micro ring resonator system. Here, we see that for two identical rings of round-trip time $T$, that the inter-MRR coupling introduces a splitting of the resonances independent of the free-spectral range of the rings. For this plot, we used critical coupling where the round-trip loss in one MRR is equal to the “reflection” coefficient off of the bus waveguide back into the MRR. Losses and coupling parameters for this plot were chosen to illustrate the structure of the transmission spectrum. By utilizing this splitting, we can use small rings with a large FSR (and therefore large finesse and cavity enhancement factor) whilst keeping the spacing between resonant frequencies of interest in the microwave band to facilitate transduction. Note: if we were to include optical backscattering, this would amount to including a second splitting on top of the already once split peaks (resulting in quadruplets). In assuming the optical backscatter is negligible, we assume this second splitting is too small to be resolved, relative to the linewidth.
III.3.1 Inter-MRR coupling constant
@@
which is precisely the form describing the system under study here. From this, we may understand that the evanescent optical coupling $J$ between MRRs determines the frequency splitting between the symmetric and antisymmetric supermodes used in this transduction scheme. Tuning this evanescent coupling amounts to changing the spacing between the MRRs, as described in the next subsection.
III.3.2 waveguide coupling constant
The waveguide coupling constant $\kappa_{ex,2}$ describes the fraction of evanescent light surrounding the bus waveguide that couples into the MRR next to it and vise versa. To adjust this coupling, one can either change the distance between the bus waveguide and the MRR, or else change the shape of the MRR to change the length over which this interaction between modes occurs.
In general, the optical waveguide coupling is dependent on the wavelength of the light being coupled. If we consider two parallel waveguides close enough to one another that the evanescent field surrounding one waveguide overlaps with the other, we can express the light propagating in the waveguides in a new basis. Instead of separate modes interacting with one another, we can consider superpositions of modes (i.e., supermodes) that are independent of one another. Light initially in one waveguide, is in an even superposition of the symmetric and antisymmetic supermodes, but these modes in general have different mode profiles, and as a result, experience different effective indices of refraction $(n_{\text{eff}}^{(sym)},n_{\text{eff}}^{(asym)})$ throughout the waveguides. Because of this accumulating phase lag between supermodes, the light overall oscillates between waveguides with a characteristic coupling beat length $L_{c}$ [30]:
$L_{c}\approx\frac{\lambda}{2(n_{\text{eff}}^{(sym)}-n_{\text{eff}}^{(asym)})}$ (29)
@@
In the transducer system considered here, we can understand that if we are considering GHz-scale detunings from a telecom-band wavelength, this is a correspondingly small change in the coupling beat length $L_{c}$, and therefore in the coupling constant. For the small detunings considered in this paper, we may take the waveguide optical coupling to be constant, but still something that may be set at the time of manufacture.
IV Discussion: Optimizing the transduction efficiency
@@
Using the equations coupling the different frequency modes in the transducer (51), one can derive an overall input-output relation for the transducer, as shown in Section B.5:
$\tilde{a}_{\text{out}}(\omega)$ $=G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}(\omega)\tilde{c}_{\text{in}}(\omega)+...$ (30a)
$G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}$ $=\frac{\sqrt{\kappa_{ex,2}}\sqrt{\gamma_{ex}}\chi_{01}\chi_{02}\chi_{m}\bar{G}\bar{a}_{1}J}{1+\bar{G}^{2}|\bar{a}_{1}|^{2}\chi_{01}\chi_{m}+J^{2}\chi_{01}\chi_{02}}$ (30b)
@@
Starting with the system parameters in Ref. [16], we can reproduce their efficiency plot as a function of driving laser power, as shown in Fig. 4 (shallow solid blue curve).
Where it is not straightforward from inspection to see which values of the system parameters will yield a maximum transduction efficiency (3), we can make this analysis much simpler by expressing (3) in terms of cooperativity parameters and relative coupling efficiency parameters (also called extraction efficiencies in [16]). Indeed, expressing the transduction efficiency in terms of cooperativities yields simplified transduction efficiency formulas used in the quantum information science community [31, 32, 33, 34, 4, 35, 36]).
The optomechanical cooperativity parameter $\mathcal{C}_{O\!M}$ will be given by the product of two ratios; the ratio of the optomechanical coupling constant $\bar{G}\bar{a}_{1}$ over the optical damping constant $\frac{\kappa_{1}}{2}$, and the ratio of the optomechanical coupling constant $\bar{G}\bar{a}_{1}$ over the mechanical damping constant $\frac{\gamma_{m}}{2}$:
$\mathcal{C}_{O\!M}\equiv\frac{4\bar{G}^{2}|\bar{a}_{1}|^{2}}{\kappa_{1}\gamma_{m}}$ (31)
As a product of ratios of the coupling between two principal modes over their loss/damping constants, cooperativities much greater than unity place us in the regime where the time scale of conversion between the two principal modes is much faster than the time scale of scattering/loss among these modes. Indeed, the square root of cooperativity is the ratio of coupling rate between modes, over the (geometric) mean loss rate per mode.
For simplicity of analysis, we can generalize cooperativity parameters to cooperativity functions (as was done in [16]) over frequency using the appropriate susceptibilities:
$\mathcal{C}_{O\!M}(\omega)=\bar{G}^{2}|\bar{a}_{1}|^{2}\chi_{01}(\omega)\chi_{m}(\omega)$ (32)
where the original cooperativity parameter is recovered on resonance.
In addition to the optomechanical coupling, we define the inter-ring optical cooperativity $\mathcal{C}_{12}$ as the cooperativity between the respective optical modes in each ring (i.e., between $\hat{a}_{1}$ and $\hat{a}_{2}$) as the corresponding product of ratios of coupling over loss:
$\mathcal{C}_{12}$ $\equiv\frac{4J^{2}}{\kappa_{1}\kappa_{2}}$ (33a)
$\mathcal{C}_{12}(\omega)$ $=J^{2}\chi_{01}(\omega)\chi_{02}(\omega)$ (33b)
With these cooperativity functions, the expression for the transduction efficiency (3) greatly simplifies:
$|\!G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}\!|^{2}\!\!=\!\frac{\kappa_{ex,2}\!\chi_{02}(\omega)\gamma_{ex}\chi_{m}\!(\omega)}{4}\frac{4\mathcal{C}_{O\!M}(\omega)\mathcal{C}_{12}(\omega)}{(1\!+\!\mathcal{C}_{O\!M}(\omega)\!+\!\mathcal{C}_{12}(\omega))^{2}}$ (34)
On resonance, we can simplify the prefactors in the first term of the transduction efficiency as a product of relative coupling efficiencies $\mathscr{F}_{2}$ and $\mathscr{F}_{m}$ (also called extraction efficiencies as they represent the fraction of outcoupling not from intrinsic losses):
$\frac{\kappa_{ex,2}\chi_{02}(\omega)}{2}$ $\equiv\mathscr{F}_{2}(\omega)\rightarrow\frac{\kappa_{ex,2}}{\kappa_{2}}$ (35a)
$\frac{\gamma_{ex}\chi_{m}(\omega)}{2}$ $\equiv\mathscr{F}_{m}(\omega)\rightarrow\frac{\gamma_{ex}}{\gamma_{m}},$ (35b)
yielding a transduction efficiency formula on resonance made of independent parameters that is also straightforward to optimize:
$|\!G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}\!|^{2}\!\!\rightarrow\!\mathscr{F}_{2}\mathscr{F}_{m}\frac{4\mathcal{C}_{O\!M}\mathcal{C}_{12}}{(1\!+\!\mathcal{C}_{O\!M}\!+\!\mathcal{C}_{12})^{2}}.$ (36)
The portion of the transduction efficiency expression excluding $\mathscr{F}_{2}$ and $\mathscr{F}_{m}$ is known as the internal conversion efficiency of the transducer.
Note: The parameters $(\mathscr{F}_{2},\mathscr{F}_{m},\mathcal{C}_{O\!M},\mathcal{C}_{12})$ are not wholly independent functions, as for example $\mathscr{F}_{m}$ contains some factors shared by $\mathcal{C}_{O\!M}$. However, they are independently tuneable, since there are enough system parameters that one may arbitrarily tune any one of $(\mathscr{F}_{2},\mathscr{F}_{m},\mathcal{C}_{O\!M},\mathcal{C}_{12})$ while holding the other three constant. Although this efficiency formula and its analysis are given in various forms in multiple references [16, 34, 33, 32, 31], our treatment gives rigorous optimization conditions to obtain a maximal transduction efficiency for intermediate levels of cooperativity.
IV.1 Tuning the couplings
IV.1.1 Straightforward optimization in terms of cooperativities
First, we note that the pump power appears only in the optomechanical coupling, and consequently, only in the optomechanical cooperativity $\mathcal{C}_{O\!M}$. This implies that one may reduce the required input pump power for optimum transduction by increasing the single-photon optomechanical coupling $G$. For the transducer in [16], $G$ is given in Table 2 as $2\pi\times 400$ Hz, but much higher (beyond three orders of magnitude) optomechanical couplings are achievable in similar piezo-optomechanical systems as seen in [37, 38, 39, 40, 41, 42].
Since all four-variables in the transduction efficiency expression (equation (36)) can be varied independently, it is straightforward to show  [^5] When all other variables are held constant, the critical value of $\mathcal{C}_{O\!M}$ that maximizes the transduction efficiency may be found by analyzing the zeroes of the derivative of the transduction efficiency with respect to $\mathcal{C}_{O\!M}$ (as is done in the calculus of smooth functions). that for any fixed set of values $(\mathscr{F}_{2},\mathscr{F}_{m},\mathcal{C}_{12})$, the value of $\mathcal{C}_{O\!M}$ that maximizes the transduction efficiency is where $\mathcal{C}_{O\!M}=\mathcal{C}_{12}+1$. This accounts for the single maximum seen in the transduction efficiency as a function of pump power plotted in Fig. 4. For large cooperativities, this also reproduces the impedance-matching rule of thumb of letting $\mathcal{C}_{O\!M}\approx\mathcal{C}_{12}$ to optimize transduction seen in the literature [34, 33, 32, 31].
Moreover, the fact that the transduction efficiency peaks at a finite power shows that, all other factors constant, optimizing the transduction efficiency is more subtle than simply increasing the coupling coefficients and decreasing the losses. For a full mathematical treatment optimizing the efficiency of a general quantum transducer employing an arbitrary number of intermediate modes, we recommend the reference [44].
In terms of (intra-ring) pump photon number $|\bar{a}_{1}|^{2}$, the critical pump power $|\bar{a}_{1,critical}|^{2}$ to obtain the maximizing value of $\mathcal{C}_{OM}$ is given by the relation:
$|\bar{a}_{1,critical}|^{2}=\frac{\gamma_{m}}{4\bar{G}^{2}}\left(\frac{4J^{2}}{\kappa_{2}}+\kappa_{1}\right)$ (37)
yielding the maximized efficiency:
$|\!G_{\tilde{a}_{\text{out}},\tilde{c}_{\text{in}}}\!|^{2}\!\!\rightarrow\!\mathscr{F}_{2}\mathscr{F}_{m}\left(\frac{\mathcal{C}_{12}}{\mathcal{C}_{12}+1}\right).$ (38)
With this partially optimized transduction efficiency expressed as a product of independent monotonic functions over the respective domains of independent variables, further maximization is a straightforward selection of whatever system parameters maximize $\mathscr{F}_{2}$, $\mathscr{F}_{m}$, and $\mathcal{C}_{12})$. Moreover, we can directly plot the maximum transduction efficiency as a function of tuning the electromechanical coupling $g_{EM}$, and the waveguide coupling constant $\kappa_{ex,2}$, as shown in Figure 5.
Figure 5: Contour plot of maximum achievable transduction efficiencies (given optimum pump power/optomechanical coupling). The axes are given in the base 10 logarithm of their frequencies (modulo $2\pi$) to smooth out the variation of the efficiency over a large range of parameters. The white dashed crosshairs give the nominal values of $g_{EM}$ and $\kappa_{ex,2}$ for the system studied in this work, corresponding to the peak of the nominal efficiency vs pump power curve in Fig. 4. The green pointed dot gives the maximum transduction efficiency when increasing $g_{EM}$ and $\kappa_{ex,2}$ each by a factor of five, as given by the peak of its corresponding curve in Fig. 4
IV.1.2 Optimal system parameters
Finding maximal values of $\mathscr{F}_{2}$, $\mathscr{F}_{m}$, and $\mathcal{C}_{12}$ can be accomplished a number of ways, such as by minimizing the intrinsic losses within $\kappa_{1}$, $\kappa_{2}$, and $\gamma_{m}$, respectively; or by maximizing the external couplings $\gamma_{ex}$ and $\kappa_{ex,2}$; or some combination of these techniques.
Where $\gamma_{ex}$ is related to the piezoelectric coupling via the relation:
$\gamma_{ex}$ $=\left|\frac{2g_{ijkmn}\sqrt{\Gamma_{ex}}}{\Gamma}\right|^{2},$ (39)
we understand that selecting materials with higher piezoelectric coefficients (all other factors constant), will increase the maximum transduction efficiency, albeit requiring a higher pump power to do so (see dotted green curve in Fig. 4). This is because increasing $\gamma_{ex}$ increases $\mathscr{F}_{m}$, which increases the maximum transduction efficiency. At the same time, increasing $\gamma_{ex}$ decreases $\mathcal{C}_{OM}$, which can be increased back to its optimum value by increasing pump power.